Inspirational Inc. is a motivational consulting business. At the end of its accounting period, October 31, 20Y2, Inspirational has assets of $763,100 and liabilities of $241,140. Using the accounting equation and considering each case independently, determine the following amounts:

a. Stockholders' equity as of October 31, 20Y2. $ 521,960

b. Stockholders' equity as of October 31, 20Y3, assuming that assets increased by $120,570 and liabilities increased by $72,490 during 20Y3. $

c. Stockholders' equity as of October 31, 20Y3, assuming that assets decreased by $60,280 and liabilities increased by $21,370 during 20Y3. $

d. Stockholders' equity as of October 31, 20Y3, assuming that assets increased by $100,730 and liabilities decreased by $38,160 during 20Y3. $

e. Net income (or net loss) during 20Y3, assuming that as of October 31, 20Y3, assets were $961,510, liabilities were $156,440, and no additional common stock was issued or dividends paid. $
Answer #1

Total Assets=Total Liabilities+Shareholders' Equity

a) Stockholders' Equity as of October 31,20Y2:-

Stockholders' Equity=Total Assets-Total Liabilities

=$763,100-$241,140=$521,960

b) Stockholders' Equity as of October 31,20Y3:-

Total Assets=763,100+120,570=883,670

Total Liabilities=241,140+72,490=313,630

Shareholders' Equity=Total Assets-Total Liabilities

=883,670-313,630=570,040

c)Stockholders' Equity as of October 31,20Y3:-

Total assets=763100-60280=702820

Total Liabilities=241140+21370=262510

Shareholders' Equity=Total Assets-Total Liabilities

=702820-262510=440310

d) Stockholders' Equity as of October 31,20Y3:-

Total assets=$763,100+100730=863830

Total Liabilities=241,140-38160=202980

Shareholders' Equity=Total Assets-Total Liabilities

=863830-202980=660850

e) Net Income:-

Total Assets=961510

Total Liabilities=156440

Shareholders' Equity=Total Assets-Total Liabilities

=961510-156440=805070

Shareholders' at the beginning=521960

This Difference is due to Net Income during the year.

So Net Income=805070-521960= 283110

"Please Rate if it helps you or let me know if anything else"